# Berlin, Germany

### v1

@@ 1 Introduction
We use intrinsically defined polynomials to model manifold-valued time-series. To this end, we utilize the extension of the ordinary de Casteljau algorithm and the resulting Bézier curves to manifolds, presented in [22] and [20]. The latter also proposes some basic properties and presents applications to motion of rigid body, space of positive definite matrices, construction of canal and developable surfaces, and curve design on implicit surfaces and polyhedra. Moreover, the mentioned work also extends the ordinary de Boor algorithm, and therewith, Bézier splines (also called composite Bézier curves) to manifolds. Note that, variational splines [6] also offer flexible models, but are computationally more costly. The work [24] characterize manifold-valued polynomials by vanishing higher-order covariant derivative. However, this approach also suffers from high computational cost (polynomials are determined by solving higher-order differential equations involving the curvature tensor). Due to evaluation via de Casteljau algorithm, Bézier polynomials offer high computational efficiency.
We employ a natural extension of the widespread least-squares regression technique to estimate the best-fitting Bézier curves that model the trajectories. We recall that smoothing, dimensional reduction, and suppressing inconsistencies and noise are advantages of regression. For a comprehensive introduction and applications, we refer to [18, 17] and [19], which employed geodesic regression (as a counter part to the ordinary linear one) and [16, 13] for the general case of manifold-valued splines. In the applications, along with the regression, we compute the coefficient of determination to examine the model’s adequacy.
Ridge regression is a popular adaptation of Tikhonov regularization to standard regression to estimate coefficients in the underlying model, especially when independent variables are highly correlated. This method has been used in many fields, including computational science and engineering and econometrics. It often provides prediction efficiency over simple linear regression with a significant number of model parameters, and helps reduce overfitting that can result from model complexity. For an overview, we refer to [15, 12] and [23]. The proposed approach in this work is a natural intrinsic extension of the ridge regression from Euclidean spaces to arbitrary Riemannian manifolds, which incorporates Mahalanobis distance and correlation matrix. It gives rise to prediction for trajectories in the underlying manifold and will be applied to forecast hurricane tracks and intensities (wind speeds).
@@
This work is organized as follows. The next section is devoted to mathematical preliminaries from Riemannian geometry, least-squares technique, the polynomial manifold-valued model and our approach, the ridge regression. In Section 3, we discuss the application of our approach to hurricanes, the modifications and numerical results.
2 Prediction for Trajectories in Riemannian Manifolds
@@
For preliminaries from Riemannian geometry, we refer to [8] and [14]. Let $M$ be a finite-dimensional Riemannian manifold. We denote its Riemannian distance by $dist$, the exponential and its local inverse, the logarithmic maps of $M$ by $\exp$ and $\log$, respectively, and set $I:=[0,1]$. Moreover, we denote the tangent space of $M$ at $x$ by $T_{x}M$, and the number of control points of the best-fitting Bézier curves (polynomials) by $n$, and set $\mathcal{M}:=M^{n}$, $\mathcal{U}:=U^{n}$, where $U$ refers to a normal convex (also called totally normal) neighbourhood in $M$.
@@ 2.1 Bézier Polynomials and Regression
Let $b_{1},\cdots,b_{n}$ be points in $U$. Applying the de Casteljau algorithm, defines a polynomial $p(t;b)$ with control points $b_{1},\cdots,b_{n}$. Here $b$ stands for $(b_{1},\cdots,b_{n})$ and $n-1$ is the degree of the polynomial $p(.;b)$. For a comprehensive introduction and some applications in data science, we refer to [16] and the recent overview [13]. Evaluation of the polynomials is, due to the use of the de Casteljau algorithm, computationally very efficient.
@@
The minimizer is the best-fitting polynomial of degree $n-1$ for data $(t_{i},y_{i})$ with $i=1,\cdots,m$. Applying a linear bijection, we may assume that $t_{1}=0$ and $t_{m}=1$. Computationally, we employ the parametrization by the control points to determine the corresponding polynomial $p(.;b^{\ast})$, where $b^{\ast}\coloneqq\operatorname{arg\,min}H$ and
@@ 2.2 Ridge Regression
We denote the Riemannian metric of $\mathcal{M}$ by $g$, and its distance, exponential and logarithmic maps by $\mathrm{d}$, $\mathrm{Exp}$ and $\mathrm{Log}$ respectively. Now, fix $\mu\in\mathcal{U}$, a positive definite selfadjoint endomorphism $S$ of $T_{\mu}\mathcal{M}$ and consider the function $F_{\mu,S}:\mathcal{U}\to\mathbb{R}$ defined by
@@
$g_{\mu}(S\log_{\mu}x,\log_{\mu}x)=g_{\mu}(W\log_{u}x,W\log_{\mu}x)$
where $W$ is a positive definite endomorphism (in the Euclidean case Tikhonov matrix) on $T_{\mu}\mathcal{M}$. Now, fix $\lambda>0$. We seek to minimize $F_{\mu,\lambda S}$ with $\mu$ the expected value, $S$ the precision, i.e., inverse of the covariance $\Sigma$ of $b$, and a suitably chosen ridge parameter $\lambda$.
@@ 2.2.1 Linearization and Gradient
Generally, due to absence of an explicit analytic solution, the regression task has to be solved numerically. To this end, we employ a Riemannian steepest descent solver [1]. In this regard, if we do not use automatic differentiation, the main challenge is the computation of the gradient of $F_{\mu,S}$. It has been shown in [4, Sec. 4.2] that $\textnormal{grad}\,H$ can be computed as the adjoint of the sum of certain Jacobi fields (in closed form, if $M$ is a symmetric space). We still need to determine the gradient of the squared Mahalanobis distance. For this purpose, we have the following result.
@@
Fix $x\in U$. Let $\gamma$ be the geodesic given by $\gamma(t)=\mathrm{Exp}_{\mu}(t\mathrm{Log}_{\mu}x)$ with $t\in I$. Let $J$ be the Jacobi field along $\gamma$ with $J(0)=0$ and $\dot{J}(0)=S\mathrm{Log}_{\mu}x$ (dot stands for derivative with respect to $t$). Then
@@
Let $k$ be the function defined by $k(x)=\mathrm{Exp}_{\mu}(W\mathrm{Log}_{\mu}x)$ on $U$, and denote the tangent map of $k$ evaluated at $x$ by $d_{x}k$. Then
@@
We remark that if $M$ is embedded in some Euclidean space, than the gradient can also be obtained by projecting the Euclidean one to the tangent space at the point of the evaluation.
@@ 2.2.3 Parameter Optimization and Forecasting
Now, fix $\lambda$ and initial samples $y_{1},\cdots,y_{i-1}$ of a discrete trajectory not in the prior. We aim to predict $y_{i}$. Our initial prediction reads $\hat{y}_{i}:=p(t_{i};b_{i}^{\ast})$, where $b_{i}^{\ast}$ denotes the minimizer of $F_{\mu,\lambda S}$. A common issue is that, in general, for coarse sampling as well as for long-term forecasting ($t_{i}-t_{i-1}$ much larger than the preceding time step), the error $d(\hat{y}_{i},y_{i})$ can increase excessively. To account for such effects and fine tuning, we slightly extend our ridge regression by adding a simple and computationally fast averaging step as follows. We replace $\hat{y}_{i}$ with $\exp_{y_{i-1}}\hat{y}_{i}$. Note that for $\alpha=1$, $\hat{y}_{i}$ remains unchanged. Iteration over $i$ yields the initial prediction for the whole trajectory. The initial and final values of $i$ are application dependent.
@@ 3.1 Dataset
We verify the effectiveness of the proposed framework by applying it to the Atlantic hurricane data from the HURDAT 2 database provided by the U.S. National Oceanic and Atmospheric Administration, publicly available on https://www.nhc.noaa.gov/data/. The data comprise measurements of latitude, longitude, and wind speed on a 6 hours base. The sample size (number of points that make up a track) varies from 13 to 96. Minimum of observed wind speeds is $15$ knots. Fig. 1 illustrates this dataset depicting the 2021 tracks together with their intensities.
@@ 3.2 Experiments and Discussion
In our experiments, following the verification rules [^1] Only tropical or subtropical cyclone stages are allowed for verification. This requirement applies to both forecast and verification times. of the U.S. National Hurricane Center (NHC), we computed the average of the MAE (mean absolute error; using spherical distance for the tracks) of the 2021 Atlantic forecasts for comparison. To this end, we represented the tracks and intensities as discrete trajectories in $S^{2}$ and $\mathbb{R}$, respectively. We recall that the Riemannian exponential and logarithmic maps for $S^{2}$ read as follows.
@@
where $x,y\in S^{2}$ with $dist(x,y)<\pi$ and $v$ orthogonal to $x$ with length $\varphi$ ($\exp_{x}0=x$). We initiated each forecast using the first sample (initial value of $i$ was $2$) and choose $n=6$, resulting in average $R^{2}$ values $>0.85$ for both fitted tracks and intensities. Fig. 2 shows two example tracks with their representation in our polynomial regression model.
@@
Obviously, one can model the tracks as images of the Euclidean best-fitting polynomials in geographic coordinates (in the latitude-longitude plane) under the standard parametrization of the sphere. Our computations show that this increases errors, demonstrating the advantage of our intrinsic approach. Figure 3 shows an example with a particularly large track error, which is mainly caused by the loop. The increase in error due to the presence of a loop, in addition to metheological anomalies, is an issue known in other approaches (cf.  [2]).
@@
Of course, we did not expect results very close to those of NHC, since our approach is generic and does not include all the relevant physical parameters, while NHC uses several more elaborate physical models. The average 12 h forecast errors of the standard NHC model OCD5 (cf. [9]) and our two primary experiments regarding the Atlantic basin for the 2021 season are summarized in table  1.
Table 1: Average 12 h forecast errors of NHC/OCD5 model and proposed approach in the Atlantic basin for the 2021 (Exp1: all, Exp2: last 5) cyclones

### v2

@@ 1 Introduction
Ridge regression, a well-established extension of Tikhonov regularization, is widely used in various fields such as computational science, engineering, and econometrics to estimate model parameters, especially in scenarios with highly correlated variables. It often improves prediction accuracy over simple linear regression by mitigating overfitting that arises from complex models [15, 12, 23]. However, many modern data types, such as directional data, shapes, or trajectories on the Earth’s surface, naturally reside on non-linear, curved spaces. Ignoring this inherent geometric structure by applying Euclidean techniques may lead to inaccurate models and suboptimal predictions. In this work, we propose a novel, intrinsic extension of ridge regression from traditional Euclidean spaces to arbitrary Riemannian manifolds. This extension leverages Riemannian least-squares fitting, empirical covariance, and the Mahalanobis distance framework, enabling robust analysis and prediction of manifold-valued time-series data. A core component of our approach is representing trajectories by best-fitting intrinsic Bézier curves, constructed via a natural generalization of the classic de Casteljau algorithm to manifold setting [20, 13]. This formulation extends geodesic regression to a polynomial regression paradigm suitable for curved spaces.
Our main contributions are threefold: 1. We formulate a natural, geometry-aware and intrinsic version of ridge regression that operates directly on manifolds, allowing for effective prediction of manifold-valued trajectories. 2. We derive explicit formulas for the gradient of the resulting objective functions, facilitating efficient numerical optimization. 3. We demonstrate the versatility of our method by applying it to the forecasting of hurricane tracks and their intensities, using real-world data.
@@
This work is organized as follows. The next section is devoted to mathematical preliminaries from Riemannian geometry, followed by a section on least-squares technique, the polynomial manifold-valued model, and our approach, the ridge regression. In Section 4, we discuss the application of our approach to hurricanes, certain modifications of the proposed approach, and numerical results.
2 Preliminaries
We provide a short intuitive background on Riemannian geometry and refer to [8] and [14] for this topic. A Riemannian manifold $M$ is a smooth space where each point has a smoothly varying inner product $g_{x}$ on its tangent space $T_{x}M$ for $x\in M$. Geodesics are curves that locally minimize distance, serving as the "straight lines" of curved spaces. The Riemannian exponential map at $x$ takes a tangent vector and maps it to a point in the manifold following the geodesic emanating from $x$. The local inverse, the logarithmic map, returns the initial velocity of a geodesic connecting $x$ to another point $y$ in $M$. Distances $dist(x,y)$ are calculated as the lengths of these geodesics. Jacobi fields are vector fields along a geodesic in a Riemannian manifold that describe how nearby geodesics diverge or converge. They arise from the second variation of the energy functional or length of geodesics and reflect the curvature of the manifold. Properties and computations that depend solely on the geometry of the manifold itself, not on an embedding space, are called intrinsic.
3 Prediction for Trajectories on Riemannian Manifolds
@@
Let $M$ be a finite-dimensional Riemannian manifold. We denote its Riemannian distance by $dist$, the exponential and its local inverse, the logarithmic maps of $M$ by $\exp$ and $\log$, respectively, and set $I:=[0,1]$. Moreover, we denote the tangent space of $M$ at $x$ by $T_{x}M$, and the number of control points of the best-fitting Bézier curves (polynomials) by $n$, and set $\mathcal{M}:=M^{n}$, $\mathcal{U}:=U^{n}$, where $U$ refers to a normal convex (also called totally normal) neighbourhood in $M$.
@@ 3.1 Bézier Polynomials and Regression
Let $b_{1},\cdots,b_{n}$ be points in $U$. Application the de Casteljau algorithm, defines a polynomial $p(t;b)$ with control points $b_{1},\cdots,b_{n}$. Here $b$ stands for $(b_{1},\cdots,b_{n})$ and $n-1$ is the degree of the polynomial $p(.;b)$. For a comprehensive introduction and some applications in data science, we refer to [16] and the recent overview [13]. Evaluation of the polynomials is, due to the use of the de Casteljau algorithm, computationally very efficient.
@@
The minimizer is the best-fitting polynomial of degree $n-1$ for data $(t_{i},y_{i})$ with $i=1,\cdots,m$. Applying a linear bijection, we may assume that $t_{1}=0$ and $t_{m}=1$. Computationally, we employ the parametrization by the control points to determine the corresponding polynomial $p(.;b^{\ast})$, where $b^{\ast}:=\operatorname{arg\,min}H$ and
@@ 3.2 Ridge Regression
We denote the Riemannian metric of $\mathcal{M}$ by $g$, and its distance, exponential, and logarithmic maps by $\mathrm{d}$, $\mathrm{Exp}$ and $\mathrm{Log}$, respectively. Now, fix $\mu\in\mathcal{U}$, a positive definite selfadjoint endomorphism $S$ of $T_{\mu}\mathcal{M}$ and consider the function $F_{\mu,S}:\mathcal{U}\to\mathbb{R}$ defined by
@@
$g_{\mu}(S\mathrm{Log}_{\mu}x,\mathrm{Log}_{\mu}x)=g_{\mu}(W\mathrm{Log}_{m}ux,W\mathrm{Log}_{\mu}x)$
where $W$ is a positive definite endomorphism satisfying $WW^{T}=S$ (in the Euclidean case represented by the Tikhonov matrix) on $T_{\mu}\mathcal{M}$. Now, fix $\lambda>0$. We seek to minimize $F_{\mu,\lambda S}$ with $\mu$ the expected value, $S$ the precision, i.e., inverse of the covariance $\Sigma$ of $b$, and a suitably chosen ridge parameter $\lambda$.
@@ 3.2.1 Linearization and Gradient
Generally, due to absence of an explicit analytic solution, the regression task has to be solved numerically. To this end, we employ a Riemannian steepest descent solver [1]. In this regard, if we do not use automatic differentiation, the main challenge is to compute the gradient of $F_{\mu,S}$. It has been shown in [4, Sec. 4.2] that $\textnormal{grad}\,H$ can be computed as the adjoint of the sum of certain Jacobi fields (in closed form, if $M$ is a symmetric space). We still need to determine the gradient of the squared Mahalanobis distance. For this purpose, we have the following result.
@@
Fix $x\in\mathcal{U}$. Let $\gamma$ be the geodesic given by $\gamma(t)=\mathrm{Exp}_{\mu}(t\mathrm{Log}_{\mu}x)$ with $t\in I$. Let $J$ be the Jacobi field along $\gamma$ with $J(0)=0$ and $\dot{J}(0)=S\mathrm{Log}_{\mu}x$ (dot stands for derivative with respect to $t$). Then
@@
Let $k$ be the function defined by $k(x)=\mathrm{Exp}_{\mu}(W\mathrm{Log}_{\mu}x)$ on $\mathcal{U}$, and denote the tangent map of $k$ evaluated at $x$ by $d_{x}k$. Then
@@
We remark that if $M$ is embedded in some Euclidean space, then the gradient can also be obtained by projecting the Euclidean one to the tangent space at the point of the evaluation.
@@ 3.2.3 Parameter Optimization and Forecasting
Now, fix $\lambda$ and initial samples $y_{1},\cdots,y_{i-1}$ of a discrete trajectory not in the prior. Starting with $i=i_{0}$, we aim to predict $y_{i}$. Our initial prediction reads $\hat{y}_{i}:=p(t_{i};b_{i}^{\ast})$, where $b_{i}^{\ast}$ denotes the minimizer of $F_{\mu,\lambda S}$. A common issue is that, in general, for coarse sampling as well as for long-term forecasting ($t_{i}-t_{i-1}$ much larger than the preceding time step), the error $d(\hat{y}_{i},y_{i})$ can increase excessively. To account for such effects and fine tuning, we slightly extend our ridge regression by adding a simple and computationally fast averaging step as follows. We replace $\hat{y}_{i}$ with $\exp_{y_{i-1}}(\alpha\log{y_{i-1}}\hat{y}_{i})$. Note that for $\alpha=1$, $\hat{y}_{i}$ remains unchanged. Iteration over $i$ yields the initial prediction for the whole trajectory. The initial and final values of $i$ are application dependent.
@@ 4 Application: Hurricanes Forecast
In [16], we represented hurricane tracks using Bézier splines (compositions of Bézier polynomials), computed average trajectories, performed tangent principal component analysis, and presented an intensity classification using support vector machines. In the following, we apply our proposed ridge regression approach, based on Bézier polynomial representation, to hurricane forecasting.
@@ 4.1 Dataset
We verify the effectiveness of the proposed framework by applying it to the Atlantic hurricane data from the HURDAT 2 database provided by the U.S. National Oceanic and Atmospheric Administration, publicly available on https://www.nhc.noaa.gov/data/. The data comprise measurements of latitude, longitude, and wind speed on a 6 hours base. The sample size (number of points that make up a track) varies from 13 to 96. Minimum of observed wind speeds is $15$ knots. Figure 1 illustrates this dataset depicting the 2021 tracks together with their intensities.
@@ 4.2 Experiments and Discussion
In our experiments, following the verification rules of the U.S. National Hurricane Center (NHC), we computed the average of the MAE (mean absolute error; using spherical distance for the tracks) of the 2021 Atlantic forecasts for comparison. To this end, we represented the tracks and intensities as discrete trajectories in $S^{2}$ and $\mathbb{R}$, respectively. We recall that the Riemannian exponential and logarithmic maps for $S^{2}$ read as follows.
@@
where $x,y\in S^{2}$ with $dist(x,y)<\pi$ and $v$ orthogonal to $x$ with length $\varphi$ ($\exp_{x}0=x$). We initiated each forecast using the first sample (initial value of $i$ was $2$) and choose $n=6$, resulting in average $R^{2}$ values $>0.85$ for both fitted tracks and intensities. Figure 2 shows two example tracks with their representation in our polynomial regression model.
@@
Obviously, one can model the tracks as images of the Euclidean best-fitting polynomials in geographic coordinates (in the latitude-longitude plane) under the standard parametrization of the sphere. Our computations show that this increases errors, demonstrating the advantage of our intrinsic approach, which incorporates the geometry of the data. Figure 3 shows an example with a particularly large track error, which is mainly caused by the loop. The increase in error due to the presence of a loop, in addition to meteorological anomalies, is an issue known in other approaches (cf.  [2]).
@@
Table 1 summarizes the average 12 h forecast errors of the standard NHC model OCD5 (cf. [9]) and our two primary experiments regarding the Atlantic basin for the 2021. Of course, we did not expect results very close to those of NHC, since our approach is generic and does not include all the relevant physical parameters, while NHC uses several more elaborate physical models. However, our approach provides a flexible framework that can complement existing physical models, especially in scenarios where physical data is sparse, computationally intensive, or when exploring shape and morphological features. Moreover, it lays a foundation for future enhancements, such as incorporating additional physical parameters that could further improve predictive performance.
Table 1: Average 12 h forecast errors of NHC/OCD5 model and proposed approach in the Atlantic basin for the 2021 (Exp1: all, Exp2: last 5) cyclones.